Montserrat Volcano Observatory

Daily Report
Report for the period 6 pm 23 January
to 6 pm 24 January 1998

Activity at the volcano today remained similar to that of the past few days. Seismicity continues to be dominated by small to moderate sized rockfall signals with a few other types of events. Between 16:00 yesterday and 16:00 today there were 26 hybrid earthquakes, 49 rockfall signals, 1 long-period and 2 volcano-tectonic earthquakes. Two of the hybrid earthquakes preceded rockfalls.

Visibility has been very poor for most of today, but some views of the dome were possible in the afternoon. These showed that there was a lot of gas production from the summit area. In particular, there appeared to be enhanced degassing immediately after a large hybrid earthquake.

A long occupation GPS survey was started today between Harris Lookout and Dagenham. The instruments will be collected tomorrow and then results will be released in a later report. Sulphur dioxide diffusion tubes were collected from sites to the north-west of the volcano to determine the levels of this gas in the atmosphere.

Today, work has been ongoing to re-establish links between MVO in Mongo Hill and the Roche's Yard and Long Ground seismic stations on the broadband seismic network.

Dust levels at the monitoring stations in houses in Woodlands, St. Peters and Mongo Hill were low again today. Slightly raised levels are still being measured at the Catholic Primary School, but these levels are not high enough to be of great concern. However, further tests are warranted to determine the source of the dust and how it relates to activity at the site or to ash in the atmosphere.

People should remain alert and listen to Radio Montserrat for further information.

Montserrat Volcano Observatory